knights inc. has completed the purchase of new dell computers, the fair value of the equipment is 873,599. the purchase agreemenr specifies an immediate down payment of 212,000 and semiannual paymenrs of 81,569 beginning at the end of 6 months for 5 years. what is the interest rate, to the nearest percent, used in discountinf this purchase transaction?
Answer #1
Calculation of Interest rate
Present value= 873599 - 212000 = $661599
Payment= $81569
N time= 10 years
Interest=?
$661599 = $81569 *PVAF(10, r%)
$661599 / 81569 = PVAF(10, r%)
8.1109 = PVAF(10,r%)

Now , PVIFA table & check the value 8.1109 for period 10 , we will get 4 %

Thus Interest rate  4 % semiannually or 8% annually